# Moonshot: Optimizing Chain-Based Rotating Leader BFT via Optimistic Proposals

### v1

@@ V Commit Moonshot
Until now, we have measured $\lambda$ in terms of $\delta$, however, this is imprecise because $\delta$ provides no way of differentiating between the performance of protocols that exchange one type of message for another. Specifically, the pipelining technique replaces two consecutive rounds of voting for one block proposal, with one round of voting for two consecutive block proposals. This means that the commit latency of a block in the pipelined setting is proportional to the dissemination time of not only the block itself, but also its child (in the best case). More to the point, pipelining essentially exchanges the cost of disseminating a second vote for the cost of disseminating a second proposal, effectively reducing commit latency when proposals take longer to disseminate than blocks.
{boxedminipage}
[ht] Commit Moonshot can be obtained by augmenting the protocol presented in Figure 2 with following steps:
  1. 1.
    Pre-commit. Upon receiving $\mathcal{C}_{v}(B_{k})$ whilst in view $v$, if ${\sf timeout\_view}_{i}<v$, multicast $\langle\mathsf{commit},H(B_{k}),v\rangle_{i}$.
  2. 2.
    Direct Commit. Upon receiving a quorum of distinct $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$ whilst in any view, commit $B_{k}$.
@@
We characterize this behavior using the modified partially synchronous model of Blum et al. [7] and assume that small messages (in this case, votes) are delivered within $\rho$ ($\delta$ in the original formulation) time while large messages (in this case, block proposals) are delivered within $\Lambda$ time after GST. Under this model, Simple Moonshot and Pipelined Moonshot both incur $\lambda$ $=2\Lambda+\rho$.
We now present a protocol with $\lambda$ $=\Lambda+2\rho$, which we call Commit Moonshot. Accordingly, when $\rho<\Lambda$ (which we assume is typically the case in practice), this protocol provides improved commit latency over those presented previously. It obtains this result through the use of explicit commit messages and while this modification is theoretically compatible with pipelining, we present it in the non-pipelined setting to obtain our result for Theorem 2. Commit Moonshot also provides $\omega$ $=\delta$ ($\Lambda$), reorg resilience and optimistic responsiveness. Moreover, while Simple Moonshot and Pipelined Moonshot require two consecutive honest leaders to guarantee a commit after GST, Commit Moonshot requires only one.
We present the two modifications required to convert Figure 2 to Commit Moonshot in Figure 3. We briefly discuss the intuition for the safety of these modifications below, and provide complete proofs in Appendix C.
@@
Per the updated commit rule given in Figure 3, $P_{i}$ commits $B_{k}$ and all of its uncommitted ancestors upon receiving a quorum (i.e. $12f+$ 1 when $n=3f+1$) of distinct $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$ messages. This remains safe because $12f+$ 1 such messages can only exist if at least $f+1$ honest nodes do not send $\mathcal{T}_{v}$. Consequently, if any block becomes certified for $v+1$ then it must have been proposed in either an optimistic or normal proposal and thus must be a child of $B_{k}$. Otherwise, $\mathcal{TC}_{v+1}$ will contain $\mathcal{C}_{v}(B_{k})$ as its highest ranked block certificate and therefore every subsequently certified block will necessarily extend $B_{k}$.
@@ VI Implementation and Evaluation
As shown in Table I, Pipelined Moonshot and Commit Moonshot equal or surpass the theoretical performance of all prior $O(n^{2})$ CRL protocols in all considered metrics. The primary question that remains, then, is whether their increased communication complexity relative to linear protocols is justified. Accordingly, we decided to implement our protocols and evaluate them against Jolteon, a linear protocol with state-of-the-art performance in most metrics and several high-quality open-source implementations
@@
We implemented all three of our protocols by modifying the code for Jolteon available in the Narwhal-HotStuff branch of the repository [29] created by Facebook Research for evaluating Narhwal and Tusk [16]. We decoupled our implementation from Narhwal and then did the same for Jolteon so that we could compare the two consensus protocols in isolation. We replaced both the Narwhal mempool and the simulated-client process by having the leaders of each protocol create parametrically sized payloads during the block creation process. We left the TCP-based network stack mostly intact and applied the few necessary changes to both implementations to ensure that any differences in performance were solely due to the differences between the consensus protocols themselves.
Our goal was to compare the throughput and latency of our Moonshot protocols with those of Jolteon under varying conditions. Firstly, we would evaluate the trade-off between $\lambda$, $\omega$ and steady-state communication complexity by running both protocols with $f^{\prime}=0$, where $f^{\prime}$ denotes the number of actual failures in the system (i.e. $f^{\prime}\leq f=\lfloor\frac{n-1}{3}\rfloor$), under varying network and payload sizes. Secondly, we would evaluate the impact of $\tau$ and reorg resilience by running each protocol in a single network with $f^{\prime}=f$ and varying leader schedules. In doing these experiments we also naturally evaluated our protocols against each other, allowing us to reveal the trade-offs incurred by pipelining and optimistic responsiveness.
We established two metrics for throughput: Firstly, the number of blocks committed by at least $12f+$ 1 nodes in the network during a run, hereafter referred to as throughput; and secondly, the average number of bytes of committed payload data transferred per second, hereafter referred to as transfer rate. For latency, we measured the average time between the creation of a block and its commit by the $2f+1th$ node.
@@
Our networks were constructed from m5.large AWS EC2 instances running Ubuntu 20.04. Each instance had a network bandwidth of up to 10Gbps [^1] https://docs.aws.amazon.com/AWSEC2/latest/UserGuide/ec2-instance-network-bandwidth.html, 8GB of memory and Intel Xeon Platinum 8000 series processors with 2 virtual cores. The instances were evenly distributed across the us-east-1 (N. Virginia), us-west-1 (N. California), eu-north-1 (Stockholm), ap-northeast-1 (Tokyo) and ap-southeast-2 (Sydney) regions. Our implementations used ED25519 signatures and constructed certificate proofs from an array of these signatures.
@@
Figure 4: Performance Overview ($f^{\prime}=0$, $p\leq 1.8$MB)
Figure 5: Performance vs. Jolteon ($f^{\prime}=0$, No Outliers)
Figure 5: Performance vs. Jolteon ($f^{\prime}=0$, No Outliers)
Figure 6: Throughput vs Latency ($n=200$, $f^{\prime}=0$, $p\leq 9$MB)
VI-A Happy Path Evaluation
We first configured all nodes to be honest in order to evaluate the trade-off between $\lambda$, $\omega$ and steady-state communication complexity. We initially tested networks of 10, 50, 100 and 200 nodes with payload sizes ranging from empty to 1.8MB to understand how the tested protocols scale under these variables. We subsequently tested additional payload sizes in the 200 node network to discover the approximate maximum transfer rate of each protocol in this setting. We ran each configuration three times, with each run lasting five minutes, and averaged the results.
As shown in Figure 5 and Table III, all Moonshot protocols produced significantly improved throughput and latency compared to Jolteon in all tested configurations. Per Table II, the 200 node network produced significant outliers under the empty and 1.8kB payload configurations, with all three protocols exhibiting $214\%-230\%$ higher throughput and $70.5\%-71.8\%$ lower latency than Jolteon, compared to the respective maximums of $70.3\%-74.4\%$ and $55.7\%-69.0\%$ seen across all other configurations. Simple Moonshot and Pipelined Moonshot produced near-identical performance in both metrics for most configurations due to the similarity of their happy-path protocols. Conversely, although Commit Moonshot produced similar throughput to these protocols, it exhibited substantially lower latency for payloads of 18kB or more due to its explicit commit messages, clearly showing the inefficiency of pipelining when blocks are large. Generally speaking, all three Moonshot protocols produced increasingly higher throughput and relatively consistent improvements to latency compared to Jolteon as the network size increased, showing that linear communication complexity is not worthwhile if it comes at the cost of reduced $\omega$ amd $\lambda$. Finally, per Figure 6, all three protocols achieved a higher maximum transfer rate with lower latency than Jolteon in the 200 node network, with Commit Moonshot producing the best results. Overall, these results show that the happy paths of our Moonshot protocols scale well and provide meaningfully decreased latency and increased throughput compared to Jolteon under the experimental conditions, with Commit Moonshot being the most efficient option.
TABLE II: Overall Performance vs Jolteon ($f^{\prime}=0$)
@@ VI-A Happy Path Evaluation
TABLE III: Performance vs Jolteon ($f^{\prime}=0$, Outliers Removed)
@@ VI-B Fallback Path Evaluation
We subsequently further evaluated the impact of pipelining along with $\tau$, reorg resilience and optimistic responsiveness, by running all protocols with a fixed $n$, $f^{\prime}$, $p$ (i.e., block payload size) and $\Delta$ under three different fair LSO/LCO leader schedules. We chose $n=100$, $f^{\prime}=33$ and $p=0$ to maximize the impact of the quadratic steady-state complexity of our protocols without risking a repeat of the outliers seen in the $n=200$, $f^{\prime}=0$ experiments. We also chose $\Delta=500ms$, a somewhat-conservative [^2] https://www.cloudping.co/grid value for this configuration that still ensured that each protocol would make it through several iterations of the leader schedules within the five minute duration of each run. As for the leader schedules, the first ($\mathcal{B}$) had all honest nodes followed by all byzantine nodes, representing the best case for non-reorg-resilient and pipelined protocols. The second ($\mathcal{WM}$) had honest-then-byzantine leaders for $2f^{\prime}$ views, followed by honest leaders for the remaining $n-2f^{\prime}$ views, representing the worst case for reorg resilient, pipelined protocols. The third ($\mathcal{WJ}$) repeated two-honest-then-byzantine for $3f^{\prime}$ views, followed by the remaining $n-3f^{\prime}$ honest, representing the worst case for non-reorg resilient, pipelined protocols.
@@
As shown in Figures 7a and 7b, Jolteon’s performance degrades enormously in the presence of failures due to its lack of reorg resilience. This is evident by the difference in its results for $\mathcal{B}$ and $\mathcal{WJ}$, with the former producing $675\%$ higher throughput and $98.1\%$ lower latency than the latter. The pipelined nature of Simple Moonshot and Pipelined Moonshot likewise caused a significant reduction in latency between the worst ($\mathcal{WM}$) and best case ($\mathcal{B}$) leader schedules for these protocols. Simple Moonshot’s $2\Delta$ wait after a failed leader (i.e. lack of Optimistic Responsiveness) caused its performance to vary more significantly than Pipelined Moonshot, while its longer view length caused a substantial decrease in throughput.
As shown by their absence from Figure 7c, both Simple Moonshot and Pipelined Moonshot failed to improve over Jolteon under $\mathcal{WM}$. More precisely, although they respectively produced $174\%$ and $448\%$ higher throughput than Jolteon, Jolteon produced $87.3\%$ and $78.3\%$ lower latency. Both of these results were a side-effect of reorg resilience: Both Moonshot protocols committed all blocks proposed by honest leaders with Byzantine successors under this schedule, but only after a significant delay. Comparatively, Jolteon lost all such blocks due to lacking this property, with only the final block of the final honest leader in the schedule being committed with a delay. Since, Jolteon commits $n-2f^{\prime}$ out of every $n$ blocks under this schedule, its relative improvement in block commit latency should increase proportionally to $n$, while its relative throughput should similarly decrease. However, we note that in this case decreased block commit latency at the cost of decreased throughput should be considered an undesirable trade-off as it does not imply a decrease in transaction commit latency.
Finally, Commit Moonshot performed consistently well regardless of the leader election schedule due to its lack of pipelining, which denies the adversary any power to delay the commit of honest blocks. Notably, it produced $702\%$ higher throughput and $99.3\%$ ($143$x) lower latency than Jolteon under $\mathcal{WJ}$. Overall, then, Commit Moonshot produced superior performance in both the happy path and the fallback path.
@@ VII Related Work
There has been a long line of work towards designing efficient BFT SMR protocols for partially synchronous networks [12, 1, 27, 23, 10, 34, 14, 20, 21, 3, 11, 18, 16, 15, 32, 19, 31, 25, 26, 22, 13]. Our work contributes to this effort by introducing the first CRL protocols to obtain both $\omega$ $=\delta$ and $\lambda$ $=3\delta$. Our protocols further provide reorg resilience, improving their recovery time after a failed leader compared to prior chain-based works that fail to achieve this property. This is especially true of both Pipelined Moonshot and Commit Moonshot, which also have low $\tau$ and are optimistically responsive. These properties come at the cost of $O(n^{2})$ steady-state communication complexity, making our protocols less performant in this metric compared to vote-aggregator-based protocols like HotStuff, however, as shown in Section VI, this trade-off is worthwhile in many settings. We presented a brief comparison between our protocols and other recent works in Section I. We now undertake a more thorough review.
@@
Like HotStuff, many other chain-based protocols [20, 21, 18, 33, 25] have focused on minimizing communication complexity, with some achieving linearity only in their steady states and others during their view-change phases as well. Recently, some [15, 25] have even achieved amortized-linear view synchronization. In all cases though, these protocols obtain steady-state linearity through the use of a designated vote-aggregator node. As we previously observed, this naturally increases their $\lambda$, $\omega$ and $\tau$ relative to our protocols, and precludes reorg resilience when the aggregator is not the original proposer. Moreover, while most nodes incur a steady-state complexity of $O(1)$ in these protocols, the proposer must still send and the aggregator must still receive, $O(n)$ messages. This imbalance means that these protocols under-utilize the available bandwidth in the point-to-point CRL setting (in which there should be no choke-points in the network and each node should have similar capabilities).
@@
PaLa [14] is a pipelined CRL protocol with $\lambda$ $=4\delta$ and $\omega$ $=2\delta$. While this improves upon the commit latencies of linear pipelined protocols with $\omega$ $=2\delta$, like [18], PaLa achieves this result at the cost of $O(n^{2})$ communication complexity in its steady state. Accordingly, it is sub-optimal in all three properties.
@@
Apollo is the only chain-based protocol with $\omega$ $=\delta$. However, it has $\lambda$ $=(f+1)\delta$ even during failure-free executions and assumes a synchronous communication model.
@@
DAG-based consensus protocols like [16, 31, 26] and [22] focus on improving block throughput. While they naturally produce and commit more blocks over a given interval than chain-based protocols by virtue of having all nodes propose in each step, they incur $O(n^{3})$ communication in doing so. While recent protocols [22, 26] in this setting have achieved $\omega$ $=\delta$, and $\lambda$ $=3\delta$ for blocks proposed by the leader, they require at least $4\delta$ to commit blocks proposed by other nodes. Consequently, since most blocks committed by these protocols are non-leader blocks, their average commit latency is still higher than our protocols. Moreover, since these protocols use pipelining, each $\delta$ corresponds to one $\Lambda$ under our model from Section V, meaning that these latencies become even more significant relative to our protocols as block size increases.
@@
A related line of works [1, 23, 27, 3, 20] has explored optimistic commits with a latency of $2\delta$ when all nodes behave honestly or in the presence $15f-$ 1 or more nodes in the system. Additionally, Giridharan et al. [19] recently proposed BeeGees, a new pipelined chained BFT SMR protocol that is able to commit without requiring consecutive honest leaders. Investigating the integration of these optimizations into our protocols represents an interesting direction for future work.

### v2

@@ V Commit Moonshot
Until now, we have measured $\lambda$ in terms of $\delta$. However, this is imprecise because $\delta$ provides no way of differentiating between the performance of protocols that exchange one type of message for another. The pipelining technique fundamentally facilitates the removal of one or more phases from a protocol by granting another phase additional meaning. In existing pipelined consensus protocols, this technique replaces two (or more) consecutive phases of voting for one block proposal, with one phase of voting for two (or more) consecutive block proposals. This means that the commit latency of a block in the pipelined setting is proportional to the dissemination time of not only the block itself, but also its child (in the best case). More to the point, pipelining essentially exchanges the cost of disseminating additional votes for the cost of disseminating additional proposals and thus naturally increases commit latency when proposals take sufficiently longer to disseminate than votes.
Commit Moonshot can be obtained by adding the following rules to the protocol for $P_{i}$ presented in Figure 3: 1. Direct Pre-commit. Upon receiving $\mathcal{C}_{v}(B_{k})$ whilst in any view $v^{\prime}$ such that $v^{\prime}\leq v$, if ${\sf timeout\_view}_{i}<v$, multicast $\langle\mathsf{commit},H(B_{k}),v\rangle_{i}$. 2. Indirect Pre-commit. Upon receiving $\mathcal{C}_{v}(B_{k})$ whilst in any view, having multicasted a commit vote for any descendant of $B_{k}$, having ${\sf timeout\_view}_{i}<v$ and having not yet multicasted $\langle\mathsf{commit},H(B_{k}),v\rangle_{i}$, multicast $\langle\mathsf{commit},H(B_{k}),v\rangle_{i}$. 3. Alternative Direct Commit. Upon receiving a quorum of distinct $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$ whilst in any view, commit $B_{k}$.
@@
We characterize this behavior using a communication model based on the modified partially synchronous model of Blum et al. [7] in which we assume that small messages (in this case, votes) are delivered within $\rho$ time while large messages (in this case, block proposals) are delivered within $\beta$ time. Moreover, we assume that after GST $\rho=[0,\min(\beta))$ and $\beta=(\max(\rho),\Delta]$. Under this model, Simple Moonshot and Pipelined Moonshot both incur $\lambda$ $=2\beta+\rho$.
Figure 5: Explicit commit votes (pictured in green) enable Commit Moonshot to commit blocks sooner than its pipelined counterparts when block proposals (pictured in blue) take sufficiently longer to disseminate than votes.
We now present a protocol with $\lambda$ $=\beta+2\rho$, which we call Commit Moonshot. Accordingly, when $\rho<\beta$ (as in Figure 5), which we assume is typical in practice, this protocol provides improved commit latency over those previously presented by integrating an explicit pre-commit phase. Like its counterparts, Commit Moonshot also obtains $\omega$ $=\delta$ ($\beta$) and provides both reorg resilience and optimistic responsiveness. Additionally, while Simple Moonshot and Pipelined Moonshot require two consecutive honest leaders to guarantee a commit after GST, Commit Moonshot requires only one.
We present the modifications required to convert Pipelined Moonshot to Commit Moonshot in Figure 4. Since Commit Moonshot retains the rules of Pipelined Moonshot, the same liveness argument that can be made for the latter also applies to the former. However, the introduction of a secondary commit path demands additional reasoning about the safety of the protocol. We present a brief intuition to this end below and provide complete proofs in Appendix C.
@@
Per the alternative commit rule given in Figure 4, $P_{i}$ commits $B_{k}$ and all of its uncommitted ancestors upon receiving a quorum (i.e. $12f+$ 1 when $n=3f+1$) of distinct $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$ messages. This remains safe because $12f+$ 1 such messages can only exist if at least $f+1$ honest nodes do not send $\mathcal{T}_{v}$. Consequently, if any block becomes certified for $v+1$ then it must have been proposed in either an optimistic or normal proposal and thus must be a child of $B_{k}$. Otherwise, $\mathcal{TC}_{v+1}$ will contain $\mathcal{C}_{v}(B_{k})$ as its highest ranked block certificate and therefore every subsequently certified block will necessarily extend $B_{k}$.
@@ VI Implementation and Evaluation
As shown in Table I, Pipelined Moonshot and Commit Moonshot equal or surpass the theoretical performance of prior $O(n^{2})$ CRL protocols in all considered metrics. The primary question that remains, then, is whether their increased communication complexity relative to linear protocols is justified. Accordingly, we decided to implement our protocols and evaluate them against Jolteon, a linear protocol with state-of-the-art performance in most metrics and several high-quality open-source implementations.
Implementation.
@@
We implemented all three of our protocols by modifying the code for Jolteon available in the Narwhal-HotStuff branch of the repository [31] created by Facebook Research for evaluating Narhwal and Tusk [16]. We decoupled our implementation from Narhwal and did the same for Jolteon so that we could compare the two consensus protocols in isolation. We replaced both the Narwhal mempool and the simulated-client process by having the leaders of each protocol create parametrically sized payloads during the block creation process, with individual payload items being 180 bytes in size. We used ED25519 signatures and constructed certificate proofs from an array of these signatures. We left the TCP-based network stack mostly intact and applied the few necessary changes to both implementations to ensure that any differences in performance were solely due to the differences between the consensus protocols themselves.
TABLE II: Observed Latencies (in ms) between AWS regions
Destination
Source us-e-1 us-w-1 eu-n-1 ap-ne-1 ap-se-2
us-east-1 5.23 61.87 113.78 167.6 197.42
us-west-1 62.88 3.69 172.17 109.89 141.54
eu-north-1 114.09 173.31 5.48 248.67 271.68
ap-northeast-1 168.04 109.94 251.63 5.99 111.67
ap-southeast-2 199.54 146.06 272.31 112.11 4.53
Region names are abbreviated versions of the Source regions.
Setting.
@@
We chose to perform our evaluation in a setting typical of modern low-latency public blockchains such as Aptos [19] to demonstrate the efficacy of our protocols when network latency is the dominating performance factor. Accordingly, we constructed moderately-sized (up to 200 nodes) wide-area networks of nodes with high bandwidth capabilities and moderate computational capabilities. Specifically, we distributed the nodes evenly across the us-east-1 (N. Virginia), us-west-1 (N. California), eu-north-1 (Stockholm), ap-northeast-1 (Tokyo) and ap-southeast-2 (Sydney) AWS regions, with each node being allocated its own m5.large EC2 instance and connected to every other node via a separate point-to-point link. Each instance ran Ubuntu 20.04 and had a network bandwidth of up to 10Gbps [^1] https://docs.aws.amazon.com/AWSEC2/latest/UserGuide/ec2-instance-network-bandwidth.html, 8GB of memory and Intel Xeon Platinum 8000 series processors with 2 virtual cores. Table II reports the typical (90th percentile) latencies observed between these regions around the time of our experiments.
Variables and Metrics.
[…]
@@
TABLE III: Performance vs Jolteon ($f^{\prime}=0$)
@@
TABLE IV: Performance vs Jolteon ($f^{\prime}=0$, Outliers Removed)
@@
Figure 6: Performance Overview ($f^{\prime}=0$, $p\leq 1.8$MB). Key trends: (1) Throughput approximately halves and latency roughly doubles for every order of magnitude increase in $p$. (2) Performance in both metrics decreases for all protocols as the network size increases. (3) Our protocols perform similarly in terms of throughput; Commit Moonshot achieves increasingly better latency as $p$ increases. (4) Our protocols outperform Jolteon in both metrics.
VI-A Happy Path Evaluation ($f^{\prime}=0$)
We initially tested networks of 10, 50, 100 and 200 honest nodes with block payload sizes ranging from empty to 1.8MB to understand how the tested protocols scale under an increasing communication load. Figure 6 reports the results of these experiments. We subsequently tested additional payload sizes in the 200 node network to discover the approximate maximum transfer rate of each protocol in this setting, which can be seen in Figure 8.
As shown in Figure 6, Figure 7 and Table IV, all Moonshot protocols produced notably higher throughput than Jolteon in all tested configurations due to the more frequent block production afforded by their reduced $\omega$. Likewise, the reduction in $\lambda$ achieved by multicasting votes (in conjunction with optimistic proposals, in the case of the pipelined protocols) also caused them to produce substantially decreased latency compared to Jolteon across all configurations. The 200 node network produced significant outliers under the empty and 1.8kB payload configurations, with all three protocols exhibiting about thrice the throughput and a quarter of the latency of Jolteon, compared to the approximately $50\%$ increase in throughput and $40\%-50\%$ reduction in latency seen on average across all other configurations. Simple Moonshot and Pipelined Moonshot produced near-identical performance in both metrics for most configurations due to the similarity of their happy-path protocols. Conversely, although Commit Moonshot produced similar throughput to these protocols, it exhibited substantially reduced latency for payloads above 18kB due to its explicit commit messages, clearly showing the inefficiency of pipelining when blocks are large. Generally speaking, all three Moonshot protocols produced increasingly higher throughput and relatively consistent improvements to latency compared to Jolteon as the network size increased, showing that obtaining linear communication complexity is counter-productive in WANs of this scale if it comes at the cost of sequentializing network operations (i.e., reducing $\omega$ and $\lambda$). Finally, per Figure 8, all three Moonshot protocols achieved a higher maximum transfer rate with lower latency than Jolteon in the 200 node network, with Commit Moonshot producing the best results. Overall, these results show that the happy paths of our Moonshot protocols scale well and provide meaningfully decreased latency and increased throughput compared to Jolteon under the experimental conditions, with Commit Moonshot being the most efficient option.
Figure 7: Performance vs. Jolteon ($f^{\prime}=0$, No Outliers)
Figure 8: Throughput vs Latency ($n=200$, $f^{\prime}=0$, $p\leq 9$MB)
@@ VI-B Fallback Path Evaluation
We subsequently further evaluated the impact of pipelining along with $\tau$, reorg resilience and optimistic responsiveness, by running all protocols with a fixed $n$, $f^{\prime}$, $p$ (i.e., block payload size) and $\Delta$ under three different fair LSO/LCO leader schedules. We chose $n=100$, $f^{\prime}=33$ and $p=0$ to maximize the impact of the quadratic steady-state complexity of our protocols without risking a repeat of the outliers seen in the $n=200$, $f^{\prime}=0$ experiments. We also chose $\Delta=500ms$, a somewhat-conservative value (per Table II) that still ensured that each protocol would make it through several iterations of the leader schedules within the five minute duration of each run. As for the leader schedules, the first ($\mathcal{B}$) had all honest nodes followed by all byzantine nodes, representing the best case for non-reorg-resilient and pipelined protocols. The second ($\mathcal{WM}$) had honest-then-byzantine leaders for $2f^{\prime}$ views, followed by honest leaders for the remaining $n-2f^{\prime}$ views, representing the worst case for reorg resilient, pipelined protocols. The third ($\mathcal{WJ}$) repeated two-honest-then-byzantine for $3f^{\prime}$ views, followed by the remaining $n-3f^{\prime}$ honest, representing the worst case for non-reorg resilient, pipelined protocols.
@@
As shown in Figures 9a and 9b, Jolteon’s performance degrades enormously in the presence of failures due to its lack of reorg resilience. This is evident by the difference in its results for $\mathcal{B}$ and $\mathcal{WJ}$, with the former producing approximately seven times higher throughput and fifty times lower latency than the latter. The pipelined nature of Simple Moonshot and Pipelined Moonshot likewise caused a significant reduction in latency between the worst ($\mathcal{WM}$) and best case ($\mathcal{B}$) leader schedules for these protocols. Simple Moonshot’s $2\Delta$ wait after a failed leader (i.e. lack of Optimistic Responsiveness) caused its performance to vary more significantly than Pipelined Moonshot, while its longer view length caused a substantial decrease in throughput.
As shown by their absence from Figure 9c, both Simple Moonshot and Pipelined Moonshot failed to improve over Jolteon under $\mathcal{WM}$. More precisely, although they both produced a several-fold increase in throughput compared to Jolteon, Jolteon produced much lower latency. Both of these results were a side-effect of reorg resilience: Both Moonshot protocols committed all blocks proposed by honest leaders with Byzantine successors under this schedule, but only after a significant delay. Comparatively, Jolteon lost all such blocks due to lacking this property, with only the block of the final honest leader in the schedule being committed with a delay. Since Jolteon commits $n-2f^{\prime}$ out of every $n$ blocks under this schedule, its relative improvement in block commit latency should increase proportionally to $n$, while its relative throughput should similarly decrease. However, we note that in this case reduced block commit latency at the cost of decreased throughput should be considered an undesirable trade-off as it does not imply a reduction in transaction commit latency.
Finally, Commit Moonshot performed consistently well regardless of the leader schedule due to its explicit pre-commit phase, which denies the adversary any power to delay the commit of honest blocks. Notably, as shown in Figure 9c, it produced around eight times higher throughput and more than two orders of magnitude lower latency than Jolteon under $\mathcal{WJ}$. Overall, then, Commit Moonshot produced superior performance in both the happy path and in the presence of failures, making it a prime candidate for application in modern low-latency public blockchains.
@@ VII Related Work
There has been a long line of work towards designing efficient BFT SMR protocols for partially synchronous networks [12, 1, 29, 25, 10, 37, 14, 22, 23, 3, 11, 20, 16, 15, 34, 21, 33, 27, 28, 24, 13]. Our work contributes to this effort by introducing the first CRL protocols to obtain both $\omega$ $=\delta$ and $\lambda$ $=3\delta$. Our protocols further provide reorg resilience, improving their recovery time after a failed leader compared to prior chain-based works that fail to achieve this property. This is especially true of both Pipelined Moonshot and Commit Moonshot, which also have low $\tau$ and are optimistically responsive. These properties come at the cost of $O(n^{2})$ steady-state communication complexity, making our protocols less performant in this metric compared to vote-aggregator-based protocols like HotStuff. However, as shown in Section VI, this trade-off is worthwhile in many settings. We presented a brief comparison between our protocols and other recent works in Section I. We now undertake a more thorough review.
@@
Like HotStuff, many other chain-based protocols [22, 23, 20, 35, 27] have focused on minimising communication complexity, with some achieving linearity only in their steady states and others during their view-change phases as well. Recently, some [15, 27] have even achieved amortized-linear view synchronization. In all cases though, these protocols obtain steady-state linearity through the use of a designated vote-aggregator node. As we previously observed, this naturally increases their $\lambda$, $\omega$ and $\tau$ relative to our protocols, and precludes reorg resilience when the aggregator is not the original proposer. Moreover, while most nodes incur a steady-state complexity of $O(1)$ in these protocols, the proposer must still send and the aggregator must still receive, $O(n)$ messages. This imbalance means that these protocols under-utilize the available bandwidth in the point-to-point CRL setting (in which there should be no choke-points in the network and each node should have similar capabilities).
@@
PaLa [14] is a pipelined CRL protocol with $\lambda$ $=4\delta$ and $\omega$ $=2\delta$. While this improves upon the commit latencies of linear pipelined protocols with $\omega$ $=2\delta$, like [20], PaLa achieves this result at the cost of $O(n^{2})$ communication complexity in its steady state. Accordingly, PaLa is sub-optimal in all three properties.
@@
Apollo obtains $\omega$ $=\delta$ at the cost of a $\lambda$ $=(f+1)\delta$ and assuming a synchronous communication model.
@@
DAG-based consensus protocols like [33, 34, 28] and [24] focus on improving block throughput. While they naturally produce and commit more blocks over a given interval than chain-based protocols by virtue of having all nodes propose in each step, they incur $O(n^{3})$ communication in doing so. While recent protocols [24, 28] in this setting have achieved $\omega$ $=\delta$, and $\lambda$ $=3\delta$ for blocks proposed by the leader, they require at least $4\delta$ to commit blocks proposed by other nodes. Consequently, since most blocks committed by these protocols are non-leader blocks, their average block commit latency is still higher than our protocols. Moreover, since these protocols use pipelining, each $\delta$ corresponds to one $\beta$ under our model from Section V, meaning that these latencies become even more significant relative to our protocols as blocks become larger.
@@
Moonshot may be further optimized by applying insights from other works. For example, a related line of works [1, 25, 29, 3, 22] achieve $\lambda=2\delta$ via optimistic commits when $n\geq 5f-1$. Similarly, works such as [36] and [17] have leveraged trusted execution environments to limit the power of the adversary, enabling consensus when $n\geq 2f+1$. Giridharan et al. also recently proposed BeeGees [21], a pipelined CRL protocol that is able to commit without requiring consecutive honest leaders. Defining variants of Moonshot that leverage these optimizations represents an interesting direction for future work.